OpenStudy (anonymous):

The area of a regular octagon is 35 cm². What is the area of a regular octagon with sides three times as long? (1 point)315 cm² 175 cm² 225 cm² 105 cm²

OpenStudy (anonymous):

I just multiply the original area by 3 right??

OpenStudy (anonymous):

hmmm, im not sure if that is quite correct. personally i would find the sides of the octagon from the area equation and then redo the problem. Area of a octagon = side^2(2+2(sqrt(2))) 35cm^2=side^2(2+2(sqrt(2)) 35/(2+2(sqrt2)=side^2 7.25=side^2 sqrt(7.25)=side=2.69

OpenStudy (anonymous):

i would then multiply each side by 3 and plug that into the equation: 2.69*3=8.07 area=side^2(2+(2sqrt(2)) area=8.07^2(2+(2sqrt(2) area=315cm^2

OpenStudy (anonymous):

Why multiply by 3?

OpenStudy (anonymous):

because each side is 3 times as long

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so that is why cant multiply the total area by 3, because it says that each side is 3times as long, not the total area is 3times as big

OpenStudy (anonymous):

Oh ok...THANK YOU!!

OpenStudy (anonymous):

no problem